Turkey Incorporated began constructing a building on the 1st of January and completed construction on the 31st of December, of the same year. Costs on the construction during this one-year period were as follows: March 1 $3,200 August 1 2,280 Dec 31 4,000 On January 1, Turkey borrowed $1,600 to finance the construction. They signed a 5-year, 12% note when they borrowed the money. Additional debt Turkey owed during the entire year included a 10%, 3-year, $1,000 note and an 11%, 4-year, $1,500 note. Round intermediate calculations to four decimals (e.g. 0.359782 rounds to 36.98%) and final answers to whole dollars. What are the weighted average accumulated expenditures?
Answer #1

ДА В | D | Date Expenditures Capitalization period Weight Mar. 1 3,200 10/12 83.3333% Aug. 1 2.280 5/12 41.6667% Dec.31 4,000

Working note:

Date Expenditures Weight Mar. 1 Aug. 1 Dec.31 Total 3200 2280 4000 9480 Capitalization period 10/12 5/12 0/12 =10/12 =5/12 =0